OpenStudy (billyjean):

Find f^-1 for the function f(x)=^3sqrt x-2 +8. A)f^-1(x)=(x-8)^3+2 B)f^-1(x)=(x+8)^3+2 C)f^-1(x)=^3sqrt x-8 +2 D)f^-1(x)=(x-8)^3-2

OpenStudy (3mar):

Well, I am here.

OpenStudy (3mar):

The original function is ?? It is not clearly typed!

OpenStudy (3mar):

One minute please

OpenStudy (3mar):

I am sorry for late!

OpenStudy (3mar):

If your problem looks like that: \[y=\sqrt[3]{x-2}+8\], we will proceed as follows;;;

OpenStudy (3mar):

Firstly, we will solve for x; i.e. separate x-terms in one side and the other terms at the other side. \[y=\sqrt[3]{x-2}+8\] \[y-8=\sqrt[3]{x-2}\] \[(y-8)^3=x-2\] \[(y-8)^3+2=x\] \[x=(y-8)^3+2\] Secondly: swap x and y; just exchange their positions \[(x-8)^3+2=y\] \[y=(x-8)^3+2\] \[f^{-1}(x)=(x-8)^3+2\] So your choice would be the first one (A)

OpenStudy (3mar):

Hope that helps

OpenStudy (phi):

what is the question ?

OpenStudy (billyjean):

i dont know if 3mar did it right or not

OpenStudy (phi):

does this f(x)=^3sqrt x-2 +8. A)f^-1(x)=(x-8)^3+2 mean \[ f(x)= \sqrt[3]{x-2}+8 \] ?

OpenStudy (billyjean):

yes

OpenStudy (phi):

3mar shows the steps. you first add -8 to both sides then "cube" both sides then add +2 to both sides you get (y-8)^3 + 2 = x then swap letters y= (x-8)^3 + 2 or, if you like \[ f^{-1}(x) = (x-8)^3 + 2 \]